OpenStudy (anonymous):

A farmer will make a rectangular pen with 100 feet of fence using part of his barn for one side of the pen. What is the largest area he can enclose?

OpenStudy (anonymous):

100/4=25*25=625

OpenStudy (anonymous):

I know the answer, but i need the work. Answer is 1250

OpenStudy (anonymous):

is that the actual answer?

OpenStudy (anonymous):

The perimeter: 2(x + y) = 100 + x The area: x * y = area 2x + 2y = 100 + x 2y = 100 + x - 2x y = 50 - 0.5x x * (50 - 0.5x) = area -0,5x^2 + 50x = area First derivative: -x + 50 Solve for x: x = 50 So y must be: 2x + 2y = 100 + x 2 * 50 + 2 * y = 100 + 50 100 + 2y = 150 2y = 50 y = 25 Test drive: x * y = 50 * 25 = 1250

OpenStudy (anonymous):

The first derivative is used for finding maximum point(s).

OpenStudy (anonymous):

Thanks :D